OpenStudy (anonymous):

find the function f(x) whose second derivaive is f"(x)=12x-sin(x)+1 and satisfies f(0)=-2 and f'(0)=2

OpenStudy (anonymous):

first anti derivative is \(f'(x)=6x^2+\cos(x)+x+c\) find \(c\) by replacing \(x\) by 0 and setting the result equal to 2

OpenStudy (anonymous):

so c is going to be 2?

OpenStudy (anonymous):

no i think \(c=1\)

OpenStudy (anonymous):

\(f'(0)=\cos(0)+c=1+c=2\) so \(c=1\)

OpenStudy (anonymous):

2x^3 + sinx + 0.5x^2 + x - 2

OpenStudy (anonymous):

how did you got that @logical reason?

OpenStudy (anonymous):

Alright... As satellite said the first derivative would be 6x^2 + cosx + x +1 Take integral again and you get 2x^3 + sinx + 0.5x^2 +x +c Substitute x=0 and you get f(0) = c = -2 Therefore answer is 2x^3 + sinx + 0.5x^2 +x - 2